Use the accompanying regression constants and coefficients to complete parts​ (a) through​ (e) below.

A. Regression constant a= 1.1 Regression coefficient b= 0.7

B. Constant= 9.8 Coefficient= -0.5

C. Constant= 2.2 Coefficient= .1

D. Constant= 9.4 Coefficient= -0.7

​(a) Determine the linear prediction rule for predicting criterion variable Y from predictor variable X using row​ (a-d)

(e) Make a single graph​ (with values from 0 to 10 on both​ axes) showing all the regression​ lines, labeling each by its letter
Answer #1

a) A. a=1l, b = 0 7 ei : Linear prediction rule 1.1 +0.7a y -0.5 B. a 9.8, b .. Linear prediction rule es y = 9.80 0.52 021 C59 A - 9 = 1 + 0 1 2 3 0 น 9 2 146 18:/, B. 9 - 98-03 24 1 0 8 78 10 18 (G 9:8 C. 9 = 2 2 4 0 1/2 0 3 76 27 | 82 9) 2:2 D 2 =